OpenStudy (anonymous):

Help!! will give medal Finish the proof that ∠AED≅∠ACB Given:BC≅DE Statements Reasons 1)BC≅DE 1)Given 2)∠BAC≅∠DAE 2)??????? 3)CB is perpendicular to AE 3)Given ED is perpendicular to AC 4) def of perpendicular 4)∠ABC and ∠ADE are rt ∠ lines 5)∠ABC≅∠ADE 5)Right angle thm 6)????????????? 6)??????????? 7))∠AED≅∠ACB 7)??????????

OpenStudy (anonymous):



OpenStudy (anonymous):

@heather040200 Sorry but it doesn't look like BAC = DAE for 2). Are you sure that is the case or is the diagram a bit off?

OpenStudy (anonymous):

yes I am sure @genius12 that one is given to me

OpenStudy (anonymous):

the diagram could be off I had to draw it

OpenStudy (anonymous):

@stgreen @Hero @robtobey

OpenStudy (anonymous):



OpenStudy (anonymous):

????

OpenStudy (anonymous):

Let <DAE=X <ADE=90 <AED=90-X (1) <ABC=90 <<ACB=90-X (2) From (1) & (2) <AED=<ACD

OpenStudy (anonymous):

CORRECTION <A E D=<A C B

OpenStudy (anonymous):

Ok but I need to pretty much fill in the blanks......the reasons should be postulates or theorems

OpenStudy (anonymous):

@Preetha

OpenStudy (anonymous):

@magbak

OpenStudy (magbak):

I am working on it

OpenStudy (anonymous):

k thanks :)

OpenStudy (magbak):

Are their any options.

OpenStudy (anonymous):

<BAC=<EAD (COMMON)

OpenStudy (anonymous):

sadly no I am just given that I just don't kno what postates or theorems to use

OpenStudy (magbak):

I know the answers but I am a little rusty with my theorems and postulates and axioms .

OpenStudy (anonymous):

triangles are similar

OpenStudy (anonymous):

OK

OpenStudy (magbak):

No he is wrong.

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

@adilalvi

OpenStudy (magbak):

2) they are the same angle. just writen in a different way. Sorry but that is all that I know.

OpenStudy (magbak):

Got to go hope I helped bye.

OpenStudy (anonymous):

@jhonyy9 @amistre64

OpenStudy (anonymous):

@RadEn